9 نقاط Find the particular solution to the : differential equation y +(+ x)cot x = 0 when y = 1 and - = 0 ع إضافة ملف
Answer #1

y + lyt yu) en(th) + Cl (1 6 + 1 17 ) giu): tca=1 C2 = 1 + 1 + ln (1/2) + ci( Eurz + J16-ne 10 where Li= (1+) ta en (Sina) +